# CoLLEGe: Concept Embedding Generation for Large Language Models

### v1

@@ 4 Datasets for Training CoLLEGe
Because CoLLEGe is designed to learn a single new token per sequence, and the LLM is frozen, training is highly sensitive to data quality, both for the support and query sequences. Additionally, three forms of mismatch between support and query sequences are important to guard against: language, contextual meaning, and knowledge mismatch. The first case is mostly self-explanatory, non-English support sequences for an English query sequence cause difficulties in training. Contextual meaning mismatch was particularly important to avoid when training with the Pile (Gao et al., 2020), whose examples are drawn from a variety of sources. Creating support and query sequences from WikiText, as in HiCE, often implicitly controls for contextual meaning (all examples are from one sources (Wikipedia), and support and query sequences for a word are often unintentionally drawn from the same article and thus share contextual meaning). Likewise, knowledge mismatch is more prominent when training with the Pile, since it contains more diverse sources. If one, or many, support sequences are more confusing than the query sequence, this can destabilize training.
Source Num. Examples
Pile-CC 79,606
Books3 51,850
BooksCorpus2 3,436
Table 1: The top Pile subsets represented in our dataset.
@@
To build a set of support sequences for each query sequence, we first split all examples into individual sentences, and matched each query sequence with sentences that use the same new word. We removed sentences that appear in the query sequence, examples with a large number of newline characters (these often were article titles, tables of contents, or lists), and examples with fewer than 15 words. In earlier experiments, we found it helpful to embed each query sequence and candidate set of support sequences with DistilRoBERTa (Sanh et al., 2019), extract the contextual embedding of the new word from both the query sequence and each candidate, and score candidates by the cosine similarity of the embeddings, in order to reduce contextual meaning mismatch. Later analysis revealed that the majority of these low-quality or mixed-meaning examples came from FreeLaw, USPTO, PubMed, or NIH ExPorter, which often have specialized meaning and formatting. Excluding those from the samples made the filtering process much more straightforward. Table 1 summarizes the top subsets from the Pile represented in our dataset.
@@
Method 1-Shot 2-Shot 3-Shot 4-Shot Def + 1-Shot Def + 2-Shot Def + 3-Shot Def + 4-Shot
@@
Table 2: Accuracy in percentage on the GRE Verbal Reasoning task, averaged over 10 trials. In each trial, a new set of $K$ example sentences are sampled and used for the $K$-shot task. For Token Tuning, we use LR = 1e-3, as that gave the best performance.
@@ 5 Experiments
In this section, we show experimental results on four different evaluation tasks that we designed: GRE verbal reasoning, definition generation, and slang identification. Note that CoLLEGe is a task-general concept embedding generation network, and all of the evaluations are performed zero-shot, without further training or fine-tuning, just like pretrained LLMs. In the following, we first discuss implementation and training details, then describe the baseline methods. Afterwards, we present the core results in Subsections 5.1-LABEL:sec:relation.
@@
As our pretrained MLM model for the Sequence Encoder, we use RoBERTa-Large (Liu et al., 2019), and apply a trainable Transformer Encoder layer to encode the RoBERTa sequence embeddings. These embeddings are aggregated using mean-pooling to produce a single embedding per sequence, which is further mean-pooled into our Concept Embedding. We use a pretrained LLaMA-2 7B model (Touvron et al., 2023) as the pretrained autoregressive language model in all our experiments.
During training, we provide 1 support sequence for each query sequence and generalize to $K>1$ during testing. We train our model for 28000 steps at batch size 32, with a learning rate of 1e-3, a linear learning rate schedule with warmup, and using the AdamW optimizer (Loshchilov & Hutter, 2019) with weight decay = 0.1 and default beta values. We experimented with different beta values during training, but found they had little effect. During training we clip gradients to a norm of 1.0. We save checkpoints based on both LM loss on the test set as well as cross entropy loss on new tokens in the test set. The final model checkpoint is selected based on its GRE score.
@@
In order to evaluate the effectivness of our method, evaluate against baselines from prior work on new concept learnign as well as prompting and gradient descent tuning. More details on implementation for the baselines can be found in Appendix C.
@@
  • Prompting uses randomly initialized new token embeddings and includes the support sequences in the prompt. It is a strong baseline with direct context access. Since prompting allows the LM to reason over the tokens in context, it can be combined with our embedding generation approach (see Appendix D for additional experiments). However, prompting original sentences can also make the context window too long and distracting.
@@ 5.1 GRE Verbal Reasoning
Using actual GRE practice questions from a Kaplan GRE prep book (Kaplan, 2019), we design a task where a language model has to either select the top or top-2 choices for a sequence with blanks. Examples for each of these questions are provided in Table 7 in Appendix A. Questions were hand-annotated from the Kaplan book and details about the cleaning process can be found in Appendix B. We produce a high-quality selection of 43 GRE Verbal Reasoning problems with a single blank (i.e. not multi-part). On an version of this task without new tokens, a pretrained LLaMa-2 7B scores 75%, which serves as an upper bound on our potential performance.
@@ Results:
Results are reported in Table 2, with additional results for “Prompting+…” reported in Table 8 in Appendix D.1. Token tuning does not seem to help much, and even sometimes hurts performance. Prompting by including the definition of each term alongside the example sentences improves performance for the baselines the most, but CoLLEGe significantly outperforms by between 4% and 19%. Model performance also increases with more examples, showing effective utilization of multiple example sentences. By contrast, more examples can sometimes harm Prompting, and Prompting+CoLLEGe, performance, likely due to the additional in-context examples distracting the LLaMA model. We note that, although our model outperforms the baselines with and without support sentences provided in-context, it performs better without them. We hypothesize that not including instruction-tuning data during the pretraining process for the new embedding reduces performance when prompted.
@@
The eerie creak of the attic door, coupled with the flickering candlelight, was enough to give anyone the <nonce>.
a feeling of unease, usually in the stomach, caused by anxiety or fear.
feelings of uneasiness
willies
Intrigued by holistic therapies, she found herself lying on a soft mat as the therapist applied <nonce> to various points on her body to alleviate her chronic migraines.
a substance that is used to heal or soothe a part of the body.
treatment of symptoms by applying pressure with the fingers to specific pressure points on the body
acupressure
Nestled in the far corner of the bustling newsroom, the diligent <nonce> worked tirelessly, transcribing reporter’s notes into clean, easy-to-read articles.
a person who writes or edits for a newspaper, magazine, or other publication.
someone employed to make written copies of documents and manuscripts
copyist
The delicate <nonce> sprouted from the forest floor, adding a touch of alien beauty to the woodland scene.
a plant that resembles a mushroom.
a fungus composed of several apothecia that look like elongated rabbit ears;
Wynnea  americana
Table 3: Definitions generated with CoLLEGe, using the prompt “The word |<nonce>| is defined as”. Each definition is generated using the single example sentence shown. None of the example sentences are provided in-context to the model.
@@ 5.2 Definition Generation
To probe how well our model understands a new word, we prompt the LLM to generate a definition for the word given a few example sentences, as shown in Figure 1.
Dataset:
To construct the dataset for evaluation, we selected 954 words from WordNet (Miller, 1994). We then prompt GPT-4 (OpenAI, 2023) to generate an example sentence for the word using the prompt: Give me a unique, descriptive sentence using the word “[WORD]” without defining it or making it obvious what the word means. Without the latter half of the prompt, many generated examples rephrased the definition. Examples are generated at temperature = 0.8. Since both our model and the baselines continue to generate text, we select the first sentence generated as the definition for scoring.
@@ Results:
Generation examples are reported in Table 3. The generated embeddings often capture high- and low-level semantic details of the example sentences. Sometimes this is fairly precise, for example the generated definition for willies is exactly correct and similarly with copyist. CoLLEGe is also able to identify Wynnea americana as a mushroom. Even when the definition is not quite right, it may capture general features of the concept correctly, and may reflect the limited information contained in the example sentence. The definition for acupressure, for example, is not exactly correct but a very good inference based on the example provided. Additional generated definitions, including those generated using more than one example sentence, are shown in Table 11 in Appendix E.1. We also show side-by-side comparisons with the baselines (without prompting) in Table 12 in Appendix E.2. Some failure cases are described in Appendix E.3.
@@
In order to evaluate the quality of generated definitions, we compare a definition generated from our model to one generated from a baseline model as well as to a ground truth definition. For comparison between models, we simulate a head-to-head competition and compute the ELO score (Elo, 1978) of each model. Specifically, for each example in the task dataset, we choose a $k$-shot setting and sample a baseline at random. We then compare the definition generated by CoLLEGe with the one generated by the baseline. Based on the result—win, lose, or tie—we update the ELO score, starting with an initial score of 1000 for all models. To choose a winner in each “round”, we use the Outlines package [^4] https://github.com/outlines-dev/outlines, we ask GPT-3.5 to select which definition is best for the word in question or if they are tied. The order of the choices (both generated definitions and “tie”) are randomized. We compute ELO separately for Table 9. We also compare the generated definitions with ground truth definitions for each word. Using the ground truth definition as the reference and the generated definition as the candidate, we compute the BERTScore F1 and report average values for our model as well as each baseline. In both quantitative evaluations, CoLLEGe outperforms the baselines. Qualitatively, only Prompting produces generated definitions that are somewhat competitive. Definitions generated by the other baselines are often incoherent, generating repetitive text or unrelated words and characters.
Model BERTScore F1 ELO
TT-1 0.752 $\pm$ 0.081 980.78
$\pm$
18.48
TT-2 0.752 $\pm$ 0.081 978.49
$\pm$
18.42
HiCE 0.767 $\pm$ 0.022 975.64
$\pm$
7.86
Additive 0.801 $\pm$ 0.023 967.22
$\pm$
8.50
Prompting 0.825 $\pm$ 0.028 1032.28
$\pm$
28.27
CoLLEGe 0.848 $\pm$ 0.023 1065.57
$\pm$
24.01
Table 4: Results for the definition generation task. We compare the model generated definitions with a reference definition generated by GPT-4 using BERTScore. Additionally, we simulate random challenges between CoLLEGe and each baseline and compute and ELO rating. For Token Tuning, we report results for LR = 3e-4, as that yielded the best results.
@@
Table 5: Accuracy in percentage on the Twitter Slang task.
Slang Term Definition Example Tweet
rizz
The ability to confidently approach people and talk to them, in a more romantic or flirty way.
Imagine having so little rizz that even the AI girlfriend rejects you. Just complete negative game.
hits different
When something is significantly better than usual or is way better under certain circumstances.
getting called pretty in person just hits different. people be making my day.
gorpcore
A fashion style that is similar to that of
hiking/wilderness/utility tech wear.
@@ 5.3 Twitter Slang
We hand-curate a set of 80 recent slang terms as well as their definitions. Alongside each term is a list of up to 8 high-quality example Tweets which use the term in an informative way. For this hand-curated set, example tweets are predominantly from 2022 and 2023. Definitions are taken from UrbanDictionary, Dictionary.com’s Pop Culture and Slang sections, the recent American Dialect Society meeting, Bark, and Wiktionary. To supplement these, we then sample 120 additional slang terms from UrbanDictionary and the Online Slang Dictionary. We select example tweets from the Twitter archive, using the downloading and processing pipeline from Hu et al. (2022). Some examples from the hand-crafted set are shown in Table 6. More information about the filtering process for this dataset can be found in Appendix B and links to the websites and curation details can be found Appendix F.
To evaluate the different models on this task, we select the true slang term, its example tweets, and its true definition. We then select 3 different incorrect slang terms and their examples. We score the log probability of the true definition conditioned on each set of examples. The highest probability is selected as the “choice”. If it corresponds to the correct combination definition for the slang term, that is counted as a correct choice, otherwise not. We score the model based on its accuracy across the whole set of slang terms.
@@ Results:
6 Conclusion and Discussion
In this paper we present CoLLEGe, a few-shot learning framework for new concept acquisition and knowledge augmentation for pretrained LLMs. We model our meta-learning approach on the original pretraining task by sampling few-shot learning episodes directly from language model pretraining datasets (e.g. the Pile) and use next-word prediction, the pretraining objective, as our primary meta-learning objective.
We find that training CoLLEGe with a single support sequence for each query sequence generalizes well to multiple support sequences at test time, and that the generated embeddings contain rich and task-general semantic information. To thoroughly evaluate the quality of these embeddings, we propose three challenging tasks including verbal reasoning, definition generation, and slang identification. We find that the generated embeddings transfer directly to these tasks zero-shot, without additional finetuning or training, and are particularly useful for more complex reasoning tasks.
While CoLLEGe achieves the best performance in all of the benchmarks, we summarize a few limitations in our current framework. First, the generated embeddings sometimes miss precise details in the examples, instead encoding higher level semantic information. This can be seen in some incorrect generated definitions, where general features of the unknown concept are correctly inferred even though specific details are missed. Second, we find the averaging mechanism cannot fully achieve parity with pretrained embeddings, even with more support sequences. For example, our best score on the few-shot GRE Verbal Reasoning task is still lower than the 75% accuracy LLaMA achieves on the task without new tokens. Additionally, Token Tuning may be more useful when we initialize the new token embedding with the embedding generated by CoLLEGe, but we do not investigate that in this paper.
Our work points to a number of future research directions. In the short term, work needs to be done to investigate different data mixes for training CoLLEGe and understand the effect of data sources on the generated embeddings (e.g. inclusion/exclusion of instruction-tuning data, non-English text, code data, etc.). More broadly, CoLLEGe only scratches the surface of exciting research in concept learning. This research is a first step in an exciting direction for future research: online continual concept acquisition performed jointly with pretraining—incrementally identifying and compressing new concepts from an online stream of sequential experience. Furthermore, this approach can be expanded towards grouping of composite novel concepts for more flexible and hierarchical organization of knowledge.
Acknowledgment

### v2

@@ 4 Datasets for Training CoLLEGe
Because CoLLEGe is designed to learn a single new token per sequence, and the LLM is frozen, training is highly sensitive to data quality, both for the support and query sequences. Additionally, three forms of mismatch between support and query sequences are important to guard against: language, contextual meaning, and knowledge mismatch. The first case is mostly self-explanatory, non-English support sequences for an English query sequence cause difficulties in training. Contextual meaning mismatch was particularly important to avoid when training with the Pile (Gao et al., 2020), whose examples are drawn from a variety of sources. Creating support and query sequences from WikiText, as in HiCE, often implicitly controls for contextual meaning (all examples are from one source (Wikipedia), and support and query sequences for a word are often unintentionally drawn from the same article and thus share contextual meaning). Likewise, knowledge mismatch is more prominent when training with the Pile, since it contains more diverse sources. If one, or many, support sequences are more confusing than the query sequence, this can destabilize training.
@@
To build a set of support sequences for each query sequence, we first split all examples into individual sentences, and matched each query sequence with sentences that use the same new word. We removed sentences that appear in the query sequence, examples with a large number of newline characters (these often were article titles, tables of contents, or lists), and examples with fewer than 15 words. Earlier experiments training with the Pile showed that some subsets had many low-quality or mixed-meaning examples. Excluding those subsets made the filtering process more straightforward. Table 6 summarizes the top subsets from the Pile represented in our dataset.
@@
Method 1-Shot 2-Shot 3-Shot 4-Shot D+1-Shot D+2-Shot D+3-Shot D+4-Shot
@@
Table 1: Accuracy in percentage on the GRE Verbal Reasoning task. For Token Tuning, we use LR = 1e-3, as that gave the best performance.
@@ 5 Experiments
In this section, we show experimental results on four different evaluation tasks that we designed: GRE verbal reasoning, definition generation, and slang identification. Note that CoLLEGe is a task-general concept embedding generation network, and all of the evaluations are performed zero-shot, without further training or fine-tuning, just like pretrained LLMs. In the following, we first discuss implementation and training details, then describe the baseline methods. Afterwards, we present the core results in Subsections 5.1-5.3.
@@
As our pretrained MLM model for the Sequence Encoder, we use RoBERTa-Large (Liu et al., 2019), and apply a trainable Transformer Encoder layer to encode the RoBERTa sequence embeddings. These embeddings are aggregated using mean-pooling to produce a single embedding per sequence, which is further mean-pooled into our Concept Embedding. We use a pretrained LLaMA-2 7B model (Touvron et al., 2023) as the pretrained autoregressive language model in all our experiments. Further implementation details can be found in Appendix B.
@@
In order to evaluate the effectivness of our method, evaluate against baselines from prior work on new concept learnign as well as prompting and gradient descent tuning. More details on implementation for the baselines can be found in Appendix E.
@@
  • Prompting uses randomly initialized new token embeddings and includes the support sequences in the prompt. It is a strong baseline with direct context access. Since prompting allows the LM to reason over the tokens in context, it can be combined with our embedding generation approach (see Appendix F for additional experiments). However, prompting original sentences can also make the context window too long and distracting.
Model BERTScore F1 ROUGE–L ELO
TT-1 75.2 $\pm$ 8.1 7.67 $\pm$ 0.1 980.78 $\pm$ 18.5
TT-2 75.2 $\pm$ 8.1 7.64 $\pm$ 0.1 978.49 $\pm$ 18.4
HiCE 76.7 $\pm$ 2.2 7.98 $\pm$ 0.0 975.64 $\pm$ 7.9
Additive 80.1 $\pm$ 2.3 11.66 $\pm$ 0.1 967.22 $\pm$ 8.5
Prompting 82.5 $\pm$ 2.8 15.54 $\pm$ 0.1 1032.28 $\pm$ 28.3
CoLLEGe 84.8 $\pm$ 2.3 17.81 $\pm$ 0.1 1065.57 $\pm$ 24.0
Table 2: Results when evaluating definition generation using the CoLLEGe-DefGen dataset. We compare the model generated definitions with a reference definition using BERTScore and also compute an ELO score for head-to-head comparison. For Token Tuning, we use LR = 3e-4, as that performed best. Definitions are generated 1-, 2-, and 3-shot and scores are averaged.
@@ 5.1 GRE Verbal Reasoning
Using actual GRE practice questions from a Kaplan GRE prep book (Kaplan, 2019), we design a task where a language model has to either select the top or top-2 choices for a sequence with blanks. Examples for each of these questions are provided in Table 7 in Appendix C. Questions were hand-annotated from the Kaplan book and details about the cleaning process can be found in Appendix D. We produce a high-quality selection of 44 GRE Verbal Reasoning problems with a single blank (i.e. not multi-part). No partial credit is given for selecting one of the two answers for a “top-2” question and chance accuracy is 8.1%. On a version of this task with the real word forms, not new tokens, a pretrained LLaMa-2 7B scores 75%, which serves as an upper bound on our potential performance.
@@ Results:
Results are reported in Table 1, with additional results for “Prompting+…” reported in Table 8 in Appendix F.1. Token tuning does not seem to help much, and even sometimes hurts performance. Model performance also increases with more examples, showing effective utilization of multiple example sentences. By contrast, more examples can sometimes harm Prompting performance, likely due to the additional in-context examples distracting the LLaMA model. We ablate the use of negative examples, knowledge distillation, and both the $L_{\text{cos}}$ and $L_{\text{mse}}$ components the knowledge distillation loss. CoLLEGe outperforms each ablation except losing to one in two columns (D+2/3-shot). Notably, neither negative examples nor knowledge distillation is that effective on its own, but when combined boost performance significantly.
@@
The eerie creak of the attic door, coupled with the flickering candlelight, was enough to give anyone the <nonce>. a feeling of unease, usually in the stomach, caused by anxiety or fear. feelings of uneasiness willies
Intrigued by holistic therapies, she found herself lying on a soft mat as the therapist applied <nonce> to various points on her body to alleviate her chronic migraines. a substance that is used to heal or soothe a part of the body. treatment of symptoms by applying pressure with the fingers to specific pressure points on the body acupressure
Nestled in the far corner of the bustling newsroom, the diligent <nonce> worked tirelessly, transcribing reporter’s notes into clean, easy-to-read articles. a person who writes or edits for a newspaper, magazine, or other publication. someone employed to make written copies of documents and manuscripts copyist
The delicate <nonce> sprouted from the forest floor, adding a touch of alien beauty to the woodland scene. a plant that resembles a mushroom. a fungus composed of several apothecia that look like elongated rabbit ears; Wynnea  americana
Table 3: Definitions generated with CoLLEGe, using the prompt “The word <nonce> is defined as”. Each definition is generated using the single example sentence shown. None of the example sentences are provided in-context to the model.
@@ 5.2 Definition Generation
To probe how well our model understands a new word, we prompt the LLM to generate a definition for the word given one or more example sentences, as shown in Figure 1.
Datasets:
We evaluate on two different datasets to test definition generation: CoLLEGe-DefGen, which we create, and the Oxford dataset used in Giulianelli et al., 2023. To construct CoLLEGe-DefGen, we selected 954 words from WordNet (Miller, 1994). We then prompt GPT-4 (OpenAI, 2023) to generate an example sentence for the word using the prompt: Give me a unique, descriptive sentence using the word “[WORD]” without defining it or making it obvious what the word means. Without the latter half of the prompt, many generated examples rephrased the definition. Examples are generated at temperature = 0.8. Since both our model and the baselines continue to generate text, we select the first sentence generated as the definition for scoring.
@@ Results:
Generation examples are reported in Table 3. The generated embeddings often capture high- and low-level semantic details of the example sentences. Sometimes this is fairly precise, for example the generated definition for willies is exactly correct and similarly with copyist. CoLLEGe is also able to identify Wynnea americana as a mushroom. Even when the definition is not quite right, it may capture general features of the concept correctly, and may reflect the limited information contained in the example sentence. The definition for acupressure, for example, is not exactly correct but a very good inference based on the example provided. Additional generated definitions, including those generated using more than one example sentence, are shown in Table 11 in Appendix G.1. We also show side-by-side comparisons with the baselines in Table 12 in Appendix G.2. Some failure cases are described in Appendix G.3.
@@
In order to evaluate the quality of generated definitions, we compare a definition generated from our model to one generated from a baseline model as well as to a ground truth definition. Definitions are generated 1-, 2-, and 3-shot and results are averaged and reported in Table 2. For comparison between models, we simulate a head-to-head competition and compute the ELO score (Elo, 1978) of each model. Specifically, for each example in the task dataset, we choose a $k$-shot setting and sample a baseline at random. We then compare the definition generated by CoLLEGe with the one generated by the baseline. Based on the result—win, lose, or tie—we update the ELO score, starting with an initial score of 1000 for all models. To choose a winner in each “round”, we use the Outlines package [^4] https://github.com/outlines-dev/outlines, we ask GPT-3.5 to select which definition is best for the word in question or if they are tied. The order of the choices (both generated definitions and “tie”) are randomized. We compute ELO separately for Table 9. We also compare the generated definitions with ground truth definitions for each word by computing the BERTScore F1 (Zhang et al., 2020) and ROUGE-L (Lin, 2004). In all three quantitative evaluations, CoLLEGe outperforms the baselines. Qualitatively, only Prompting produces generated definitions that are somewhat competitive. Definitions generated by the other baselines are often incoherent, generating repetitive text or unrelated words and characters.
CoLLEGe-DefGen Oxford
Model BERTScore F1 ROUGE-L BERTScore F1 ROUGE-L
FLAN–Base–DefInstr 84.0 12.1 83.4 15.4
FLAN-Large–DefInstr 85.0 13.2 83.6 16.6
FLAN–XL–DefInstr 83.7 11.6 83.6 14.0
CoLLEGe 83.5 16.5 83.1 15.4
Prompting + CoLLEGe 84.1 18.0 83.6 17.1
Table 4: Results for our comparison with Giulianelli et al. (2023)'s FLAN-T5 models on our CoLLEGe-DefGen dataset and the Oxford dataset. All evaluations are done 1-shot, with a single example sentence, and the target word is replaced with a new token.
We next compare CoLLEGe with the series FLAN-T5 (Chung et al., 2024) models finetuned by Giulianelli et al. (2023) to generate definitions, which we denote FLAN-Base–DefInstr, FLAN-Large–DefInstr, and FLAN-XL–DefInstr. Each of these models is prompted with an example sentence followed by the question, “What is the definition of [WORD]?”. For a fair comparison, we replace the target word with a new token when evaluating the FLAN-T5 models and restrict CoLLEGe to one example sentence, since the finetuned FLAN-T5 models are trained to only use one example as well. Finally, since the FLAN-T5 models are prompted, we include results for Prompting + CoLLEGe as well. The results are presented in Table 4. When evaluated on our dataset, both CoLLEGe and Prompting + CoLLEGe achieve much higher ROUGE-L scores than all FLAN-T5 models. For BERTScore, Prompting + CoLLEGe outperforms all but FLAN-Large, although all scores are fairly close together. On the Oxford dataset, CoLLEGe and Prompting + CoLLEGe score comparably to the FLAN-T5 models, and Prompting + CoLLEGe slightly outperforms in terms of ROUGE-L. Notably, CoLLEGe achieves this zero-shot, without any additional training on the definition generation task, while the FLAN-T5 models are specialized for this purpose.
@@
Table 5: Accuracy in percentage on the CoLLEGe-Slang benchmark dataset. CoLLEGe achieves higher accuracy on our Twitter Slang task than each of the baselines, including Prompting.
@@ 5.3 Twitter Slang
To build CoLLEGe-Slang, we first hand-curate a set of 80 recent slang terms as well as their definitions. Alongside each term is a list of up to 8 high-quality example Tweets which use the term in an informative way. For this hand-curated set, example tweets are predominantly from 2022 and 2023. To supplement these, we then sample 120 additional slang terms from UrbanDictionary and the Online Slang Dictionary. We select example tweets from the Twitter archive, using the pipeline from Hu et al. (2022). Some examples from the hand-crafted set are shown in Table 13. More information about the filtering process, data sources, and curation details for this dataset can be found in Appendix D.
To evaluate the different models on this task, we select the true slang term, its example tweets, and its true definition. We then select 3 different incorrect slang terms and their examples. We score the log probability of the true definition conditioned on each set of examples. The highest probability is selected as the “choice”. If it corresponds to the correct combination of definition and slang term, that is counted as a correct choice, otherwise not. We score the model based on its accuracy across the whole set of slang terms.
@@ Results:
6 Conclusion
In this paper we present CoLLEGe, a few-shot learning framework for new concept acquisition for pretrained LLMs. We model our meta-learning approach on the original pretraining task by sampling few-shot learning episodes from language model pretraining datasets and using next-word prediction, as our primary meta-learning objective. CoLLEGe generated embeddings contain rich and task-general semantic information, generalize to multiple challenging tasks at test time zero-shot, without additional finetuning or training, and are particularly useful for more complex reasoning tasks.
7 Limitations and Future Work
While CoLLEGe achieves the best performance in all of the benchmarks, we summarize a few limitations in our current framework. First, the generated embeddings sometimes miss precise details in the examples, instead encoding higher level semantic information. This can be seen in some incorrect generated definitions, where general features of the unknown concept are correctly inferred even though specific details are missed. Second, we find the averaging mechanism cannot fully achieve parity with pretrained embeddings, even with more support sequences.
Our work points to a number of future research directions. In the short term, work needs to be done to investigate different data mixes for training CoLLEGe. More broadly, this research is a first step in an exciting direction for future research: online continual concept acquisition performed jointly with pretraining—incrementally identifying and compressing new concepts from an online stream of sequential experience.
Acknowledgments